# Unveiling Latent Causal Rules: A Temporal Point Process Approach for Abnormal Event Explanation

### 03 Feb 2024 v1

@@ Abstract
In high-stakes systems such as healthcare, it is critical to understand the causal reasons behind unusual events, such as sudden changes in patient’s health. Unveiling the causal reasons helps with quick diagnoses and precise treatment planning. In this paper, we propose an automated method for uncovering “if-then” logic rules to explain observational events. We introduce temporal point processes to model the events of interest, and discover the set of latent rules to explain the occurrence of events. To achieve this, we employ an Expectation-Maximization (EM) algorithm. In the E-step, we calculate the likelihood of each event being explained by each discovered rule. In the M-step, we update both the rule set and model parameters to enhance the likelihood function’s lower bound. Notably, we optimize the rule set in a differential manner. Our approach demonstrates accurate performance in both discovering rules and identifying root causes. We showcase its promising results using synthetic and real healthcare datasets.
@@ 1 Introduction
However, effectively implementing abnormal event detection comes with challenges. A major obstacle is the lack of labeled abnormal event data during training. Labeling events as normal or abnormal is often expensive and labor-intensive. Consider the context of credit card transactions. When the system encounters an unusually large transaction, determining its underlying rationale can be intricate. It may indeed be a legitimate and routine transaction, or it could be indicative of fraudulent activity. Identifyineng the true reasons require thorough investigation, which is both time and resource consuming. In our paper, we address this issue by introducing a rule-based probabilistic approach for predicting and explaining events. We treat event’s explanatory reasons as hidden variables and use an EM algorithm to learn rules that explain events without explicit label information regarding anomaly. For each observed event, the triggered rules will provide insight into the most probable causal factors and determine whether the event should be flagged as abnormal or normal.
We introduce temporal point process to model events of interest in continuous time. The time-to-event is treated as a random variable characterized by the intensity function, indicating the occurrence rate of the event in real time. We formulate the intensity function as a mixture model, where each component is defined by an “if-then” rule factor. We assume that the “if” condition should occur before the “then” part, and the “if” condition also specifies the temporal order constraints that the relevant logic variables must meet, such as “if $X_{1}$ and $X_{2}$ are true and $X_{1}$ happens before $X_{2}$, then $Y$ is true.” These explanatory temporal logic rules will be discovered from data, which compete to influence when an event occurs. When an event happens, the specific rule that is triggered based on preceding events will provide an explanation for its occurrence.
@@
Notably, learning the rule set poses a formidable challenge due to its inherently combinatorial nature. It requires selecting the proper subsets of logic variables and then considering their temporal order relations to create a logic rule. This variable selection process is both time-consuming and non-differentiable. In our paper, we propose to solve a relaxed continuous optimization problem by directly modeling the logic variable selection probability in forming a rule. Specifically, each rule is generated by a reparameterized K-subset sampling with the Gumbel noise injected to make the rule sampling process differentiable [18]. We further use a padding trick to make the generated rules with various lengths not exceeding K. In the M-step, both the rule set and the model parameters can be learned end-to-end in a differentiable way.
@@
(i) We leverage the continuous-time temporal point process modeling framework and formulate the intensity function as a rule-informed mixture model. Our rule-based probabilistic model is inherently interpretable, ensuring safe and reliable event prediction and explanation.
(ii) We design an EM learning algorithm that can accommodate the latent logical reasons for each event. Our overall learning algorithm is efficient and differentiable, which can automatically discover the set of explanatory logic rules from the population data and determine the most likely logical reason for each event.
(iii) We conduct empirical evaluations of our algorithm using both synthetic and real datasets, demonstrating its strong performance in rule discovery and root cause determination.
@@ 2 Related Work
Temporal Point Process (TPP) Models
TPP models provide an elegant tool for modeling the time intervals or time-to-event in continuous time. Current research in this area primarily focuses on enhancing the flexibility of intensity functions. For example, [4] introduced the RMTPP, a neural point process model that employs a Recurrent Neural Network to model the intensity function. [12] further improved RMTPP by developing a continuous-time RNN. [23] and [21] leveraged the self-attention mechanism to capture the long-term dependencies of events. However, relying on flexible black-box models, which can excel at event prediction with sufficient training data, presents challenges. These models lack interpretability, making it difficult to understand the occurrence of events. In scenarios like abnormal event detection, the objective goes beyond merely predicting unusual events; it also involves provding causal explanations for these occurrences. The absence of interpretability renders black-box models unsuitable for high-stakes systems where transparency is critical.
The contemporary AI perspective underscores the value of crafting intrinsically interpretable models over relying solely on explaining black-box models [14]. Along this line of research, recently, [11; 10] proposed to design a rule-informed intensity function, trying to add the interoperability. However, their approaches lack the capability to provide detailed logical explanations for individual events, in contrast to our method. Additionally, their rule learning process is non-differentiable, relying on a branch and bound algorithm.
Rule Set Mining
Rule set discovery in an unsupervised fashion has long been a classic topic. Traditional approaches include Itemset Mining Methods like Apriori [1] and NEclatcloesed [2], which target frequent itemset identification without accounting for event order. Sequential Pattern Mining methods, typified by CM-SPADE [6] and VGEN [7], aim to unveil temporal relationships but they lack precision and cannot handle timestamp information. In addition to the unsupervised methods, Inductive Logic Programming (ILP) [16] provides a powerful supervised rule learning method, yet requiring balanced positive and negative examples to have a good performance. In contrast, our method leverages fine-grained temporal information to learn temporal logic rules from data and in an unsupervised manner.
@@ 3 Problem Setup
We define a set of logic variables (or predicates) as $\mathcal{X}$, where each variable $X_{u}\in\mathcal{X}$ is a boolean variable. Our target event, such as a sudden change in patient’s health or an unusual large transaction, is denoted as $Y$ and $Y\in\mathcal{X}$. We add a temporal dimension to the logic variables and each grounded logic variables becomes a sequence of spiked events, denoted as $\{X_{u}(t)\}_{t\geq 0}$, where each $X_{u}(t)\in\{0,1\}$, $\forall t$. Specifically, $X_{u}(t)$ jumps to 1 (i.e., true) at the time step when the event happens. In our problem, each data sample is a $|\mathcal{X}|$-dimensional multivariate temporal processes, denoted as $\mathcal{H}_{t}=\{X_{u}(t)\}_{u=1,\dots,|\mathcal{X}|}$.
We are interested in modeling and providing insight into the occurrence of $\{Y(t)\}_{t\geq 0}$. We treat the time intervals as random variables and the duration until event $Y$ happens is characterized by the conditional intensity function, denoted as $\lambda(t\mid\mathcal{H}_{t-})$ where $t-$ means up to $t$ but not include $t$. By definition, $\lambda(t\mid\mathcal{H}_{t-})dt=\mathbb{E}[N([t,t+dt])\mid\mathcal{H}_{t-}]$, where $N(([t,t+dt])$ denotes the number of events occurring in the interval $[t,t+dt)$. Given the occurrence time of event $Y$, i.e., $(t_{1},\dots,t_{n})$, the joint likelihood function is given by
@@
Here, to simplify the notation, we denote $p^{*}(t):=p(t\mid\mathcal{H}_{{t}-})$ and $\lambda^{*}(t):=\lambda(t\mid\mathcal{H}_{t-})$. In this paper, we will model $\lambda^{*}(t)$ using the rule-based features, through which we aim to understand the underlying logical reasons to explain the occurrence of target events. Specifically, our goals include:
(i) Discover the rule set :=$\mathcal{F}:=\{f_{1},f_{2},...,f_{H}\}$ from the entire events $\mathcal{H}_{t_{n}}$ up to $t_{n}$, with each rule having a general form
$f:\quad Y\leftarrow\left(\bigwedge\limits_{X_{u}\in\mathcal{X}_{f}}X_{u}\right)\bigwedge\left(\bigwedge\limits_{X_{u},X_{v}\in\mathcal{X}_{f}}r_{j}(X_{u},X_{v})\right)$ (2)
where $\mathcal{X}_{f}$ is the set of body predicates of rule $f$. ${r_{j}(X_{u},X_{v})}$ is a set of temporal relations between pairs of predicates and can take any of the three types, i.e., “Before”, “Equal”, and “After”, and can also be none, meaning there are no temporal order constants existing for $X_{u}$ and $X_{v}$.
@@
  • spontaneous events (i.e., cannot be explained by any logic rule or it’s meaningless to be explained);
  • can be explained by one logic rule from the discovered rule set =$\mathcal{F}=\{f_{1},f_{2},...,f_{H}\}$.
@@ 4 Model
We first describe how to construct the intensity function for $Y$ based on the logic-informed features. For a general rule as defined in Eq. (2), we can construct the rule-informed feature as
$\phi_{f}(\mathcal{H}_{t})=\prod_{X_{u}\in\mathcal{X}_{f}}X_{u}(t_{u})\cdot\prod_{X_{u},X_{v}\in\mathcal{X}_{f}}r_{j}(X_{u},X_{v})\in\{0,1\}$ (3)
where we need to check from $\mathcal{H}_{t}$ whether each $X_{u}\in\mathcal{X}_{f}$ has been once grounded as true and whether the temporal relation grounded by their times holds (for multiple events we choose the last one). The temporal relation $r_{j}(X_{u},X_{v})$ is grounded according to its type,“Before”, “Equal”, or “After” (denoted as $R_{b},R_{e},R_{a}$ respectively) by the event times as follows,
$R_{b}\left(t_{u},t_{v}\right)$ $=\mathbbm{1}\left\{t_{u}-t_{v}<-\delta\right\}$
$R_{e}\left(t_{u},t_{v}\right)$ $=\mathbbm{1}\left\{\left|t_{u}-t_{v}\right|<\delta\right\}$
$R_{a}\left(t_{u},t_{v}\right)$ $=\mathbbm{1}\left\{t_{u}-t_{v}>\delta\right\}$
where $\delta$ (can be 0) is specified as the time tolerance. If the body condition of $f$ is grounded to be true given history, then the feature $\phi_{f}(\mathcal{H}_{t})$ is 1, otherwise it is 0. A better illustration can be found in Fig. 1.
Figure 1: Illustration of the feature construction. For the logic rule: $f:Y\leftarrow X_{1}\bigwedge X_{2}\bigwedge X_{3}$, whenever the body condition becomes trues, the rule gets fired, as a result the feature $\phi_{f}$ becomes 1, and the intensity function of the head predicate is boosted by $r_{f}$.
Given the rule-informed feature defined in Eq. (3), we can build our intensity model for $\{Y(t)\}_{t\geq 0}$ as follows. For the $i$-th event in the dataset, by considering all the rules =$\mathcal{F}=\{f_{1},f_{2},...,f_{H}\}$, we model the intensity function as a mixture-of-components form,
$\lambda^{*}(t_{i}\mid\bm{z}_{i})=b_{0}+\sum_{h=1,\dots,H}\gamma_{h}z_{ih}\phi_{f_{h}}(\mathcal{H}_{t_{i}})$ (4)
where $b_{0}$ is the base term, $h$ is the index of the rule, and $\gamma_{h}>0$ denotes the impact weight of each rule $f_{h}$. The latent variable $\bm{z}_{i}=[z_{ih}]_{h=1,\dots,H}$ is a one-hot vector. When $z_{ih}=1$, it means event $i$ is due to the $h$-th rule factor, which provides logical reasons to explain its occurrence. We assume the categorical $\bm{z}_{i}$ has a distribution $\bm{\pi}=[\pi_{h}]$, which indicates the probability of a rule appearing in the population and is our learnable parameter.
In the next section, we design an EM algorithm, which aims to jointly learn the rule set and point process model parameters from data, and simultaneously perform the inference of each $\bm{z}_{i}$.
@@ 5 Learning: EM Algorithm
    In the E-step, the posterior probability over $z_{i}=[z_{ih}]_{h=0,\dots,H}$ for each event $Y^{i}$ is computed. We use $z_{i0}$ to indicate that the event cannot be explained by any discovered rule.
@@
Notably, we devise a series of techniques in EM to enhance the stability of the optimization process. The E-step has been successfully resolved in a closed-form solution. The M-step is more involved, for which we employ three-fold approximations and reparameterization to facilitate the end-to-end differentiable training.
Complete Data Likelihood
To simplify the derivation, let’s start with the $i$-th event $Y(t_{i})$ and consider its complete data likelihood. The following derivations can be easily extended to the entire event sequence.
For an occurred event $Y(t_{i})$ at time $t_{i}$, given the rule index $\bm{z}_{i}=[z_{ih}]_{h=0,\dots,H}$, we can write the likelihood for the complete data $\{(t_{i},\bm{z}_{i})\}$ as

### 19 Mar 2024 v2

@@ Abstract
In high-stakes systems such as healthcare, it is critical to understand the causal reasons behind unusual events, such as sudden changes in patient’s health. Unveiling the causal reasons helps with quick diagnoses and precise treatment planning. In this paper, we propose an automated method for uncovering “if-then” logic rules to explain observational events. We introduce temporal point processes to model the events of interest, and discover the set of latent rules to explain the occurrence of events. To achieve this goal, we employ an Expectation-Maximization (EM) algorithm. In the E-step, we calculate the posterior probability of each event being explained by each discovered rule. In the M-step, we update both the rule set and model parameters to enhance the likelihood function’s lower bound. Notably, we will optimize the rule set in a differential manner. Our approach demonstrates accurate performance in both discovering rules and identifying root causes. We showcase its promising results using synthetic and real healthcare datasets.
@@ 1 Introduction
However, effectively implementing abnormal event detection comes with challenges. A major obstacle is the lack of labeled abnormal event data during training. Labeling events as normal or abnormal is often expensive and labor-intensive. Consider the context of credit card transactions. When the system encounters an unusually large transaction, determining its underlying rationale can be intricate. It may indeed be a legitimate and routine transaction, or it could be indicative of fraudulent activity. Identifying the true reasons requires thorough investigation, which is both time- and resource-consuming. In our paper, we address this issue by introducing a rule-based probabilistic approach for predicting and explaining events. We treat the event’s explanatory reasons as hidden variables and use an EM algorithm to learn rules that explain events without explicit label information regarding anomaly. For each observed event, the triggered rules will provide insight into the most probable causal factors and determine whether the event should be flagged as abnormal or normal.
We introduce temporal point process to model events of interest in continuous time. The time-to-event is treated as a random variable characterized by the intensity function, indicating the occurrence rate of the event in real-time. We formulate the intensity function as a mixture model, where each component is defined by an “if-then” rule factor. We assume that the “if” condition should occur before the “then” part, and the “if” condition should also specify the temporal order constraints that the relevant logic variables must meet, such as “if $X_{1}$ and $X_{2}$ are true and $X_{1}$ happens before $X_{2}$, then $Y$ is true.” These explanatory temporal logic rules will be discovered from data, which compete to influence when an event occurs. When an event happens, the specific rule that is triggered based on preceding events will provide an explanation for its occurrence.
@@
Notably, learning the rule set poses a formidable challenge due to its inherently combinatorial nature. It requires selecting the proper subsets of logic variables and then considering their temporal order relations to create a logic rule. This variable selection process is both time-consuming and non-differentiable. In our paper, we propose to solve a relaxed continuous optimization problem by directly modeling the logic variable selection probability in forming a rule. Specifically, each rule is generated by a reparameterized K-subset sampling with the Gumbel noise injected to make the rule sampling process differentiable [31]. We further use a dummy predicate trick to accommodate for various rule lengths, while still being limited by $K$. In the M-step, both the rule set and the model parameters can be learned end-to-end in a differentiable way.
@@
  • We leverage the continuous-time temporal point process modeling framework and formulate the intensity function as a rule-informed mixture model. Our rule-based probabilistic model is inherently interpretable, ensuring safe and reliable event prediction and explanation.
  • We design an EM learning algorithm that can accommodate the latent logical reasons for each event. Our overall learning algorithm is efficient and differentiable, which can automatically discover the set of explanatory logic rules from the population data and determine the most likely logical reason for each event.
  • We conduct empirical evaluations of our algorithm using both synthetic and real datasets, demonstrating its strong performance in rule discovery and root cause determination.
@@ 2 Related Work
Temporal Point Process (TPP) Models TPP models provide an elegant tool for modeling the interevent time intervals or time-to-event in continuous time. Current research in this area primarily focuses on enhancing the flexibility of intensity functions. For example, Du et al. [7] introduced the RMTPP, a neural point process model that employs a Recurrent Neural Network (RNN) to model the intensity function. Mei and Eisner [23] further improved RMTPP by developing a continuous-time RNN. Zuo et al. [37] and Zhang et al. [34] leveraged the self-attention mechanism to capture the long-term dependencies of events. However, understanding how these flexible black-box models work, which excel at event prediction with sufficient training data, presents challenges. It is difficult to interpret the underlying mechanisms driving the occurrence of events from these models. In scenarios like abnormal event detection, the objective goes beyond merely predicting unusual events; it also involves providing causal explanations for these occurrences. The absence of interpretability renders black-box models unsuitable for high-stakes systems where transparency is critical.
Recently, some researchers in AI started emphasizing the importance of developing inherently interpretable models rather than solely relying on explaining black-box models [26]. Along this line of research, [19; 20] designed a rule-informed intensity function for TPPs. The highly structured intensity function, encoding logical rules as prior knowledge, enables the discovery of logic rules to explain events. Furthermore, these models can perform effectively in small data regimes. However, their approach tends to discover global explanatory logic rules from population data to maximize likelihood [20], while paying less attention to identifying the most probable explanatory logic rules at the individual event level, unlike our method. Additionally, their rule learning algorithm is non-differentiable, relying on a branch-and-bound approach to progressively add rules to the rule set.
Rule Set Mining The Discovery of explanatory rule sets from event sequences in an unsupervised manner has long been an interesting topic. Traditional approaches, such as Itemset Mining Methods like Apriori [1] and NEclatclosed [3], focus on identifying frequent event items from population data to generate association rules but overlook event order. Sequential Patterns Mining methods like CM-SPADE [10] and VGEN [11], aim to uncover temporal relationships from event sequences but often lack precision and struggle with the fine-grained recorded timestamp information. In contrast to these unsupervised methods, supervised learning methods such as Inductive Logic Programming (ILP) [29] offer a principled rule learning approach, albeit usually requiring balanced positive and negative examples for optimal performance. The rule set is discovered under the principle that all positive samples are accounted for by at least one rule, while none of the negative samples are covered by the rule set. As a comparison to all previous methods, our rule discovery method based on TPPs can naturally deal with fine-grained temporal information of events, and the learning algorithm is unsupervised, without the need for labeled positive and negative event samples.
@@ 3 Problem Setup
We define a set of logic variables (or predicates) as $\mathcal{X}$, where each variable $X_{u}\in\mathcal{X}$ is a boolean variable. Our target event, such as a sudden change in a patient’s health or an unusually large transaction, is denoted as $Y$ and we also let $Y\in\mathcal{X}$.
Adding a temporal dimension to the logic variables, we assume that each grounded logic variable from the observations is a sequence of spiked events, denoted as $\{X_{u}(t)\}_{t\geq 0}$, where each $X_{u}(t)\in\{0,1\}$, $\forall t$. Specifically, $X_{u}(t)$ jumps to 1 (i.e., true) at the time step when the event happens. In our problem, we consider the grounded logic variables as degenerated events, and each data sample observed up to time $t$ is represented as a $|\mathcal{X}|$-dimensional multivariate TPP, denoted as $\mathcal{H}_{t}=\{X_{u}(t)\}_{u=1,\dots,|\mathcal{X}|}$.
We are interested in modeling and learning logic rules from history $\mathcal{H}_{t}$ to provide insight into the occurrence of $\{Y(t)\}_{t\geq 0}$. We consider a continuous-time setting, where the distribution of the duration until event $Y$ happens is characterized by the conditional intensity function, denoted as $\lambda(t\mid\mathcal{H}_{t-})$. Here, the intensity function is conditional on history, and $t-$ means up to $t$ but not including $t$. By definition,
$\lambda(t\mid\mathcal{H}_{t-})dt=\mathbb{E}[N([t,t+dt))\mid\mathcal{H}_{t-}],$
where $N([t,t+dt))$ denotes the number of events occurring in the interval $[t,t+dt)$. Given the occurrence time of event $Y$, such as $(t_{1},\dots,t_{n})$, the joint likelihood function of these events is given by
@@
where we denote $p^{*}(t):=p(t\mid\mathcal{H}_{{t}-})$ and $\lambda^{*}(t):=\lambda(t\mid\mathcal{H}_{t-})$ to simplify the notation.
In this paper, we will model $\lambda^{*}(t)$ using the rule-informed features, and through learning the rules we aim to understand the underlying logical reasons to explain the occurrence of target events. Specifically, our goals include:
(i) Discover the rule set :=$\mathcal{F}:=\{f_{1},f_{2},...,f_{H}\}$ from the entire events $\mathcal{H}_{t_{n}}$ up to $t_{n}$, with each rule having a general Horn clause form
$f:\quad Y\leftarrow\left(\bigwedge\limits_{X_{u}\in\mathcal{X}_{f}}X_{u}\right)\bigwedge\left(\bigwedge\limits_{X_{u},X_{v}\in\mathcal{X}_{f}}R_{j}(X_{u},X_{v})\right),$ (2)
where $\mathcal{X}_{f}$ is the set of the involved body predicates in the rule $f$. ${R_{j}(X_{u},X_{v})}$ is the temporal relation between the selected predicates $X_{u}$ and $X_{v}$. $R_{j}$ can take any of the four types, i.e., “Before”, “Equal”, “After”, and “None”, specifying the temporal order constraints for the predicate pair. Notably, the temporal relations can be none, meaning there are no temporal order constraints.
@@
  • spontaneous events (i.e., cannot be explained by any logic rule or it is meaningless to be explained);
  • can be explained by one logic rule from the discovered rule set =$\mathcal{F}=\{f_{1},f_{2},...,f_{H}\}$ and our algorithm will determine the most probably one.
@@ 4 Model
We first describe how to construct the intensity function for $\{Y(t)\}_{t\geq 0}$ based on the logic-informed features. For a general rule as defined in Eq.(2), we can construct the rule-informed feature as
$\phi_{f}(\mathcal{H}_{t})=\prod_{X_{u}\in\mathcal{X}_{f}}X_{u}(t_{u})\cdot\prod_{X_{u},X_{v}\in\mathcal{X}_{f}}R_{j}(t_{u},t_{v})\in\{0,1\}$ (3)
where we need to check from $\mathcal{H}_{t}$ whether each $X_{u}\in\mathcal{X}_{f}$ has been once grounded as true at $t_{u}$ where $t_{u}<t$, and whether the temporal relation holds, which is grounded by their times (for multiple events we choose the last one). The temporal relation $R_{j}(t_{u},t_{v})$ is grounded according to its type, “Before”, “Equal”, “After”, or “None” (denoted as $R_{b},R_{e},R_{a},R_{n}$ respectively) by the event times as follows,
$R_{b}\left(t_{u},t_{v}\right)$ $=\mathbb{I}\left\{t_{u}-t_{v}<-\delta\right\}$
$R_{e}\left(t_{u},t_{v}\right)$ $=\mathbb{I}\left\{\left|t_{u}-t_{v}\right|\leq\delta\right\}$
$R_{a}\left(t_{u},t_{v}\right)$ $=\mathbb{I}\left\{t_{u}-t_{v}>\delta\right\}$
$R_{n}\left(t_{u},t_{v}\right)$ $=1$
where $\delta$ (can be 0) is specified as the time tolerance. If the body condition of $f$ is grounded to be true given history, then the feature $\phi_{f}(\mathcal{H}_{t})$ is 1; otherwise it is 0. A better illustration of how to construct the feature can be found in Fig.1.
Figure 1: Illustration of the feature construction. For a logic rule $f$ such as: $Y\leftarrow X_{1}\bigwedge X_{2}\bigwedge X_{3}$, whenever the body condition becomes true, the rule gets fired. As a result, the feature $\phi_{f}$ becomes 1, and the intensity function of the head predicate is boosted.
Given the rule-informed feature defined in Eq.(3), we can build our intensity model for $\{Y(t)\}_{t\geq 0}$ as follows. For the $i$-th event in the dataset, by first assuming all the rules =$\mathcal{F}=\{f_{1},f_{2},...,f_{H}\}$ are known, we model the intensity function as a mixture-of-components form,
$\lambda^{*}(t_{i}\mid\bm{z}_{i})=z_{i0}b_{0}+\sum_{h=1,\dots,H}\gamma_{h}z_{ih}\phi_{f_{h}}(\mathcal{H}_{t_{i}})$ (4)
where $b_{0}$ is the base term, $h$ is the index of the rule, and $\gamma_{h}>0$ denotes the impact weight of each rule $f_{h}$. The latent variable $\bm{z}_{i}=[z_{ih}]_{h=0,1,\dots,H}$ is a one-hot vector. When $z_{ih}=1$, $h\neq 0$, it means event $i$ is due to the $h$-th rule factor, which provides logical reasons to explain its occurrence. We use $z_{i0}$ to indicate that the event cannot be explained by any discovered rule. We assume that each categorical $\bm{z}_{i}$ has the same prior distribution $\bm{\pi}=[\pi_{h}]_{h=0,1,\dots,H}$, which indicates the probability of each component appearing in the population and is our learnable parameter.
In the next section, we design an EM algorithm, which aims to jointly learn the rule set and the TPP model parameters from data. Meanwhile, in the E-step, we perform the inference of each $\bm{z}_{i}$, which indicates the most probably explanatory rule for event $i$.
@@ 5 Learning: EM Algorithm
    In the E-step, the posterior probability over $\bm{z}_{i}=[z_{ih}]_{h=0,\dots,H}$ for each event $Y(t_{i})$ is computed.
@@
Notably, we devise a series of techniques in EM to enhance the stability of the optimization process. The E-step can be successfully resolved in a closed-form solution. The M-step is more involved, for which we employ three-fold approximations and reparameterization to facilitate the end-to-end differentiable training, which will be discussed later.
Complete Data Likelihood To simplify the derivation, let’s first focus on the $i$-th event $Y(t_{i})$ and consider its complete data likelihood. The following derivations can be easily extended to the entire event sequence.
For the $i$-th event $Y(t_{i})$, given the rule index $\bm{z}_{i}=[z_{ih}]_{h=0,\dots,H}$, we can write the likelihood for the complete data $\{(t_{i},\bm{z}_{i})\}$ as